OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the circumference of this circle?

OpenStudy (anonymous):

i dont know, all the information i have is here. but sadly i have no idea how to solve this

jimthompson5910 (jim_thompson5910):

Hint: C = 2*pi*r

OpenStudy (anonymous):

so 18.84?

jimthompson5910 (jim_thompson5910):

Yes approximately if you use pi = 3.14

jimthompson5910 (jim_thompson5910):

Now we use this formula s = (t/360)*C where.. s = arc length t = central angle C = circumference of entire circle

jimthompson5910 (jim_thompson5910):

Do you see where to go from here?

OpenStudy (anonymous):

kinda whats the central angle though and how would i find it

jimthompson5910 (jim_thompson5910):

That's the unknown you're solving for. It turns out that when you say "measure of an arc" you also mean "the measure of the central angle"

jimthompson5910 (jim_thompson5910):

s = (t/360)*C s = (t/360)*18.84 ... plug in the circumference you just found 4.189 = (t/360)*18.84 ... plug in the arc length given now solve for t

OpenStudy (anonymous):

ok that makes since then :)

jimthompson5910 (jim_thompson5910):

I'm glad it does